# Noise-induced synchronization in coupled quantum oscillators

### v2

@@ II.1 Steady State Solutions
It is useful to examine the long-time or stationary state solutions in which
@@
$\Delta=\frac{\left(\gamma_{12}\left(n_{1}+n_{2}+1\right)+2\gamma\sqrt{n_{1}n_{2}}\right)}{2\left(\gamma^{2}-\gamma_{12}^{2}\xi^{2}\right)}.$ (35)
@@
The off-diagonal $\tilde{\Theta}_{12}$ terms corresponds to the steady-state “flow” of quanta between the two oscillators. Interestingly, if there is a temperature bias between the two oscillators, we develop correlation since $\tilde{\Theta}_{12}\neq 0$. In fact, if we consider the time evolution of the population of oscillator 1:
$\frac{dn_{1}(t)}{dt}=\gamma(\overline{n}_{1}-n_{1}(t))-ig(\langle a_{1}^{\dagger}a_{2}\rangle+\langle a_{2}^{\dagger}a_{1}\rangle).$ (37)
This is a continuity equation since we have quanta entering from the bath (1st term) and leaving to the other oscillator (2nd term). For the stationary state, the two must be equal
$\gamma(\overline{n}_{1}-n_{1}(t))=ig(\langle a_{1}^{\dagger}a_{2}\rangle+\langle a_{2}^{\dagger}a_{1}\rangle)=J$ (38)
and
@@
gives the flux of quanta from the warmer to cooler baths. This does not depend upon correlations within the environment and only depends upon the exchange coupling term within the Hamiltonian itself.
$\Theta+i\Omega\geq 0$ (40)
where $\Omega$ is the symplectic form
II.2 Information Theory for Gaussian States
Figure 3: (a) Mutual information shared by two oscillators in the presence of a temperature bias as specified by the average occupations $n_{1}$ and $n_{2}$. (b) Mutual information shared by two oscillators at a common temperature. In each case, we set $\omega_{1}=\omega_{2}=1$, $g=1$, and $\gamma=0.1$. The vertical dashed lines correspond to the limiting conditions imposed by Eq.36 when $n_{1}\neq n_{2}$.
@@ II.2 Information Theory for Gaussian States
$\Theta_{AB}=\left[\begin{array}[]{cc}\Theta_{A}&\Gamma\\ \Gamma^{\dagger}&\Theta_{B}\end{array}\right]$ (48)
then the covariance within subspace A or B is given by the two diagonal blocks and the correlations between the two are the off-diagonal blocks ($\Gamma$). The mutual information is then simply
$)\displaystyle I_{2}(A:B)=S_{A}+S_{B}-S_{AB}=\frac{1}{2}\ln{\frac{\text{det}\Theta_{A}\text{det}\Theta_{B}}{\text{det}\Theta_{AB}}}$ (49)
where we used the Rényi-2 entropy. This mutual information is a measure of the total correlation between the two oscillators.
@@
$\begin{split}{\cal D}_{2}&=I_{2}(A:B)-{\cal J}_{2}(A|B)\\ &=\inf_{\Sigma_{B}^{\Pi}}\frac{1}{2}\ln\left(\frac{\text{det}\Theta_{B}\text{ det}\tilde{\Theta}_{A}^{\Pi}}{\text{det}\Theta_{AB}}\right).\end{split}$ (52)
In Ref. 22 we showed that one can obtain a tight lower bound on the amount of quantum correlation between subsystems by simply removing the coherences between the two subspaces within the total density matrix and computing the mutual information between the two remaining diagonal blocks. For Gaussian systems, this follows from Eq. 50 since ${\cal J}_{2}(\rho_{A|B})$ is extremized under this condition and Eq. 52 becomes equal to the total correlation.
In our previous work on coupled qubit dimers, we found that depending upon the correlation within the environment as specified by $\xi$, by and large, coupling to a common reservoir introduces quantum correlations between the spins.[21, 22] For the case at hand, we have two possible sources of correlation in our model. First, the direct coupling introduces correlation when the two oscillators are at different temperatures. Secondly, even when the direct coupling is turned off or the two oscillators are at the same temperature, correlations within the common environmental mode introduces correlations between the two oscillators. Our results are consistent with the formal results in Ref. 32 which show that the only two-mode Gaussian states with zero Gaussian quantum discord are product states, i.e. states with no correlations at all, that constitute a zero measure set. Consequently, all correlated two-mode Gaussian states have nonclassical correlations certified by a nonzero quantum discord.
@@ III Discussion
We demonstrate here that synchronization and entanglement is a universal phenomina that can arise via indirect coupling to a correlated dissipative environmental mode. Environmental correlation imposes a symmetry on the system by selectively increasing the relaxation rates for modes that have the same symmetry as the environmental correlations leaving the remaining modes effectively decoupled from the environment. This selectivity will persist in multi-mode systems, especially in system with a high degree of topological symmetery. One can use this to engineer dissipation and decoherence free domains within the total Hilbert space of the system.
@@ Acknowledgments
The authors acknowledge discussions with Prof. Carlos Silva (U. Montreal) and Andrei Piryatinski (LANL) that led to the development of this model and series of calculations. Both authors discussed the results and contributed to the final manuscript.

### v3

@@ II.1 Steady State Solutions
It is useful to examine the long-time or stationary state solutions (denoted by “tilde”) in which
@@
$\Delta=\frac{\left(\gamma_{12}\left(\overline{n}_{1}+\overline{n}_{2}+1\right)+2\gamma\sqrt{\overline{n}_{1}\overline{n}_{2}}\right)}{2\left(\gamma^{2}-\gamma_{12}^{2}\xi^{2}\right)}.$ (42)
@@
II.2 Correlation under thermal bias
The off-diagonal $\tilde{\Theta}_{12}$ terms corresponds to the steady-state “flow” of quanta between the two oscillators. Interestingly, if there is a temperature bias between the two oscillators, the system develops correlation since $\tilde{\Theta}_{12}\neq 0$. However, a temperature bias does not create synchronisation (or anti-synchronisation) since the sign of $\gamma_{12}$ in the equations of motion for the moments does not change if $\bar{n}_{1}\neq\bar{n}_{2}$. In fact, if we consider the time evolution of the population of oscillators 1 and 2 at different temperatures:
$\frac{dn_{1}(t)}{dt}=\gamma(\overline{n}_{1}-n_{1}(t))-ig(\langle a_{1}^{\dagger}a_{2}\rangle+\langle a_{2}^{\dagger}a_{1}\rangle)$ (44)
$\frac{dn_{2}(t)}{dt}=\gamma(\overline{n}_{2}-n_{2}(t))+ig(\langle a_{1}^{\dagger}a_{2}\rangle+\langle a_{2}^{\dagger}a_{1}\rangle).$ (45)
For the stationary state (in which populations do not evolve with time), the two time-derivatives must be equal and opposite in sign, and we can define a flux
@@ II.2 Correlation under thermal bias
describing the flow of quanta from the warmer to cooler baths. This does not depend upon correlations within the environment and only depends upon the exchange coupling term within the Hamiltonian and the local system/bath coupling described by $\gamma$.
II.3 Information Theory for Gaussian States
Figure 4: (a) Mutual information shared by two oscillators in the presence of a temperature bias as specified by the average occupations $\overline{n}_{1}$ and $\overline{n}_{2}$. (b) Mutual information shared by two oscillators at a common temperature. In each case, we set $\omega_{1}=\omega_{2}=1$, $g=1$, and $\gamma=0.1$. The vertical dashed lines correspond to the limiting conditions imposed by Eq.43 when $\overline{n}_{1}\neq\overline{n}_{2}$.
@@ II.3 Information Theory for Gaussian States
$\Theta=\left[\begin{array}[]{cc}\Theta_{A}&\Gamma\\ \Gamma^{\dagger}&\Theta_{B}\end{array}\right]$ (52)
such that the covariance within subspace A or B is given by the two diagonal blocks and the correlations between the two are the off-diagonal blocks ($\Gamma$). The mutual information is then simply
$)\displaystyle I_{2}(A:B)=S_{A}+S_{B}-S_{AB}=\frac{1}{2}\ln{\frac{\text{det}\Theta_{A}\text{det}\Theta_{B}}{\text{det}\Theta}}$ (53)
in terms of the Rényi-2 entropy. This mutual information is a measure of the total correlation between the two oscillators.
@@
$\begin{split}{\cal D}_{2}&=I_{2}(A:B)-{\cal J}_{2}(A|B)\\ &=\inf_{\Sigma_{B}^{\Pi}}\frac{1}{2}\ln\left(\frac{\text{det}\Theta_{B}\text{ det}\tilde{\Theta}_{A}^{\Pi}}{\text{det}\Theta}\right).\end{split}$ (56)
In Ref. 25 we showed that one can obtain a tight lower bound on the amount of quantum correlation between subsystems by simply removing the coherences between the two subspaces within the total density matrix and computing the mutual information between the two remaining diagonal blocks. For Gaussian systems, this follows from Eq. 54 since ${\cal J}_{2}(\rho_{A|B})$ is extremized under this condition and Eq. 56 becomes equal to the total correlation.
In our previous work on coupled qubit dimers, we found that depending upon the correlation within the environment as specified by $\xi$, by and large, coupling to a common reservoir introduces quantum correlations between the spins.[24, 25] For the case at hand, we have two possible sources of correlation in our model. First, the direct coupling introduces correlation when the two oscillators are at different temperatures. Secondly, even when the direct coupling is turned off or the two oscillators are at the same temperature, correlations within the common environmental mode introduces correlations between the two oscillators. Our results are consistent with the formal results in Ref. 28 which show that the only two-mode Gaussian states with zero Gaussian quantum discord are product states, i.e. states with no correlations at all, that constitute a zero measure set. Consequently, all correlated two-mode Gaussian states have nonclassical correlations certified by a nonzero quantum discord.
@@ III Discussion
We demonstrate here that synchronization and entanglement is a universal phenomina that can arise via indirect coupling to a correlated dissipative environmental mode. Environmental correlation imposes a symmetry on the system by selectively increasing the relaxation rates for modes that have the same symmetry as the environmental correlations leaving the remaining modes effectively decoupled from the environment. This selectivity will persist in multi-mode systems, especially in system with a high degree of topological symmetery. One can use this to engineer dissipation and decoherence free domains within the total Hilbert space of the system. An example of enhanced collective behavior is the recent report of quantum collective motion in a superconducting circuit optomechanical platform composed of six-equivalent oscillators in a hexagonal arrangement.[15] While the authors do not consider the influence of the environmental noise, the experiments do suggest that similar devices could be fabricated that could have controllable degrees of environmental correlations.
@@ Acknowledgments
The authors acknowledge discussions with Prof. Carlos Silva (U. Montreal) and Andrei Piryatinski (LANL), which led to the development of this model and series of calculations. Both authors discussed the results and contributed to the final manuscript.